Find the domain of the composition function f o g
F(x) = 1/ x-9 ; g(x) = 1/x





Answer by stanbon(75887)

Find the domain of the composition function f o g
F(x) = 1/ x-9 ; g(x) = 1/x
----------------------------
fog(x) = f[g(x)] = f[1/x] = 1/[(1/x)-9] = 1/[(1-9x)/x] = x/(1-9x)
--------
Domain::: ??
The dinominator cannot be zero.
Solve 1-9x = 0
9x = 1
x = 1/9
---------
Domain is "All Real Numbers except x = 1/9"
=====================================
Cheers,
Stan H.
===================